### v1

@@ 5.1 Joint LC & RV analysis with linear ephemerides
As mentioned in Section 4, we set and as free parameters under the control of a uniform prior, which implies assuming linear ephemerides. With this setup, we improved the transit depth precision of all three planets by a factor 1.4 if compared with the results of Vanderburg et al. (2019). This improvement level is consistent with having twice the number of data points with respect to the light curve analysis performed by Vanderburg et al. (2019), as well as with low TTV amplitudes (see Section 5.2).
@@
When combining the mass and radius values of the three planets, we obtain the following median estimates for the bulk planetary densities: , (), and g cm-3. We note that the RV-undetected TOI-396 c would be the least dense planet, while the densest planet is the outermost one (i.e. TOI-396 d), which constitutes a quite atypical architecture within the observed exoplanet population (e.g. Ciardi et al., 2013; Weiss et al., 2018; Mishra et al., 2023). However, this conclusion is just tentative, given the uncertainties on the mean planetary densities. Moreover, the detection level of the RV-inferred parameters of TOI-396 c is not statistically significant. We further tested whether including the MINERVA-Australis (Addison et al., 2019) RV data (30 measurements as taken from Vanderburg et al., 2019) can help detecting the elusive planet. However, it turned out that their precision level ( 6 m s-1) is not high enough to improve the characterisation of the system. In other words, the RV semi-amplitudes we obtained are consistent and indistinguishable within the statistical fluctuation with what was derived from the more precise HARPS dataset. This led us to further check that TOI-396 c indeed belongs to this system (Section 5.2) and to investigate the reason for its RV non-detection (Section 5.3).
5.2 Joint LC & RV analysis accounting for TTVs
@@
In line with Vanderburg et al. (2019), we also confirmed that the period ratio is commensurable and differs from the 5:3 ratio by less than 0.027%. As planets with orbits in, or close to, resonances are likely to show TTVs, we decided to repeat the same MCMC analysis outlined above, but enabling the transit timings of each transit event to vary to then compute the TTV amplitude with respect to the linear ephemerides model derived in Section 5.1. All jump parameters converged (). The medians of the posterior distributions of the most relevant system parameters along with the 68.3% confidence intervals are listed in Table 2. The phase-folded LCs of the three planets are shown in Figure 1, while the phase-folded RV time series are displayed in Figure 2. In particular, the middle panel of Fig. 2 shows that, after subtracting the RV signals of both planets b and d, the RV time series looks flat consistently with the non-detection of TOI-396 c.
@@
This analysis gives , , and and confirms the RV detection of TOI-396 b and TOI-396 d (at the 3.8 and 4.5-level, respectively) with and . These outcomes are consistent with the results obtained from the analysis that assumes linear ephemerides (Section 5.1).
Fig. 3 shows the three planets of the system (star symbol) along with other exoplanets whose density is more significant than the 3 level999The planet properties were downloaded from the Nasa Exoplanet Archive as of 5 June 2024: https://exoplanetarchive.ipac.caltech.edu/ (circular marker) in the mass-radius (MR) diagram. The superimposed theoretical MR models as taken from Aguichine et al. (2021, A21) and Haldemann et al. (2024) help to guide the eye, however we warn the reader of possible degeneracies occurring in the MR plane. A thorough internal structure analysis of the well characterised TOI-396 b and d is presented in Section 7.
@@
5.3 Why do not we detect TOI-396 c in the RV time series?
@@
In order to investigate this hypothesis, we performed a frequency analysis of the line profile variation diagnostics (FWHM, contrast, and skewness) and activity indicators (log R and H indexes). Figure 5 displays the time series (left column), along with the respective generalised Lomb-Scargle (GLS, Zechmeister & Kürster, 2009) periodograms (right column). We assessed the significance of the peaks detected in the power spectra by estimating their false alarm probability (FAP), i.e., the probability that noise could produce a peak with power higher than the one we found in the time series. To account for the possible presence of non-Gaussian noise in the data, we estimated the FAP using the bootstrap randomisation method (see, e.g., Murdoch et al., 1993; Kuerster et al., 1997; Hatzes, 2019). Briefly, we computed the GLS periodogram of 105 mock time series obtained by randomly shuffling the data points along with their error bars, while keeping the timestamps fixed. We defined the FAP as the fraction of those mock periodograms whose highest power exceeds the power of the real data at any frequency. In the present work, we considered a peak to be significant if its false alarm probability is FAP 0.1 %.
@@
To check whether a quasi-periodic signal compatible with 7 d is also present in the photometric data, for each TESS sector we extracted custom LCs from pixel data using lightkurve (Lightkurve Collaboration et al., 2018). In detail, we adopted the default quality bitmask and set the aperture to ‘all’, which corresponds to an aperture larger than the one used by the official SAP pipeline. In fact, larger apertures mitigate the effect of slow image drifts that could interfere with slow flux changes, such as the 6–8 d rotation period signals we aim to detect. After removing the temporal windows containing the transit events, we computed the GLS periodograms of these lightkurve-based LCs. The FAP was computed following the same bootstrap technique outlined above for the RV activity indicators. The four periodograms (Fig. 15, first column) exhibit very significant peaks at 7.7, 7.9, 7.5, and 6.8 days for TESS Sectors 3, 4, 30, and 31, respectively. Except for Sector 30, they are not the most prominent peaks, however they persist even after removing the most significant signals (Fig. 15, second column). Whereas we acknowledge that the likely rotation period of the star is close to the first harmonic of the orbital period of TESS around the Earth ( 14 days), the photometric signal at 6–8 d is significant in all the four TESS sectors and persists after pre-whitening the data. This signal is consistent with the rotation period detected in the HARPS activity indicators and inferred from log R, , and gyrochronology, suggesting it is astrophysical in nature and due to the presence of active regions carried around by stellar rotation. Assuming the log R-based d as our reference estimate, the orbital period of TOI-396 c ( d) is close to , which may explain the non-detection of planet c within the RV time series.
@@
As is not exactly equal to , we then repeated the test outlined above, but this time we injected into the original HARPS time series artificial Keplerian signals with . The values obtained by the MCMC analyses depending on the different values are reported in Table 4. The values are systematically and significantly smaller than the corresponding values, which a-posteriori supports the conclusion that the stellar rotation period is around 6–8 d. Furthermore, a planet with this orbital period would be firmly detected if its RV semi-amplitude were greater than , which is the largest RV semi-amplitude detected for this system. Instead, by injecting a Keplerian signal with and , the MCMCI analysis is able to barely detect (at 2) a planetary signal whose amplitude is half of that expected. The detection level increases when increases, however we still underestimate . These tests prove that it is difficult to retrieve planetary signals with .
In summary, we conclude that stellar activity is responsible for generating spurious RV signals whose harmonics also include the stellar rotation period. As a consequence, it is hard to reliably detect planets with orbital periods comparable to via the RV technique and Table 4 quantifies the magnitude of this effect. We note that the we obtained from the MCMCI analysis in Section 5.2 is comparable to the retrieved when inserting an artificial signal having m s-1, which let us to speculate that TOI-396 c might have 3.0 and 2.0 g cm-3.
@@ 6 Joint RV and TTV dynamical analysis
As mentioned, the anti-correlation pattern of the observed TTVs (see Fig. 14) is a typical signal of the dynamical interaction between TOI-396 b and c. However, data in our hands are not enough to currently derive meaningful planetary masses from the TTVs. Indeed, the photometric observations are clustered in two groups (about two years apart), which results in a partial coverage of the curvature of the TTVs. This prevents us from accurately mapping the full phases and super-periods of the TTV signals, hence a dynamical fit onto the transit times would lead to a high fraction of low-amplitude TTVs with short super-period or to a low fraction of high-amplitude long-period TTV signals.
@@
Despite this, we attempted to run a dynamical joint fit of the TTV and RV data set with TRADES121212https://github.com/lucaborsato/trades (Borsato et al., 2014, 2019, 2021), a Fortran-python code developed to model TTVs and RVs simultaneously along with N-body integration. We have taken and fixed the stellar mass and radius values from Table 1. We further fixed the orbital inclinations to the values in Table 2 and the longitude of the ascending nodes to 131313 We assumed the same reference system of Winn (2010) and Borsato et al. (2014) and later used in Borsato et al. (2019, 2021, 2024); Nascimbeni et al. (2023, 2024). for all the planets. We fitted the planetary masses scaled by the stellar mass (), the orbital periods (), the eccentricities () and the arguments of the pericentre () in the form (, ), and the mean longitudes (141414The mean longitude is defined as , where is the argument of the pericentre, is the longitude of the ascending node, and is the mean anomaly.). We also fitted for an RV offset () and for an RV jitter term () by adopting as step parameter, although we used the detrended RV data set as derived from Section 5.2, where a jitter term was already included in the RV errors. All fitting parameters were subject to uniform priors that account for their respective physical boundaries (see Table 5); for the eccentricities we applied a log-penalty () based on the half-Gaussian () from Van Eylen et al. (2019).
@@
The reference time for the dynamical integration of the orbital parameters was set at , that is before all available observations. We combined the quasi-global differential evolution (Storn & Price, 1997) optimisation algorithm implemented in PyDE (Parviainen et al., 2016) with the Affine Invariant MCMC Ensemble sampler (Goodman & Weare, 2010) emcee implemented by Foreman-Mackey et al. (2019, 2013). We first ran PyDE and evolved 68 different initial configurations of parameter sets for generations (number of steps for which each parameter is evolved). To perform the dynamical analysis in an MCMC fashion, we then ran emcee, assuming as starting point the outcome obtained with PyDE. We set up 68 chains for steps each. To sample the parameter space efficiently, we mixed the DEMove() and DESnookerMove() differential evolution moves151515See emcee documentation: https://emcee.readthedocs.io/en/stable/tutorials/moves/ in the proportion 80%–20% (Nelson et al., 2014; ter Braak & Vrugt, 2008). We repeated the sequence PyDE+emcee twice, with different seeds for the random number generator. The chains reached convergence according to visual inspection and statistical indicators, such as the Gelman-Rubin , the Geweke’s statistic (Geweke, 1991), and the auto-correlation function161616See e.g. Goodman & Weare (2010) and the emcee documentation at https://emcee.readthedocs.io/en/stable/tutorials/autocorr/#autocorr.. For both runs, we applied a conservative thinning factor of 100 and discarded the first of the chains (burn-in). For each run, we derived the reference outcome (hereinafter also referred to as best-fit), as the maximum-a-posteriori (MAP) parameters’ set, that is the set of parameters that maximises the log-probability171717 The log-probability is given by , where is the log-likelihood and is the log-penalty of the eccentricities. . The uncertainty of each parameter is quantified by the high-density interval (HDI) at 68.27%181818 An HDI at 68.27% is the equivalent of the percentile range of a Normal distribution that provides the 1 standard confidence interval. of its posterior distribution. For each parameter we computed a Z-score defined as , where the subscripts denote the two different runs and . It turned out that for each parameter, which allows us to merge the posterior distributions deriving from the two runs to finally compute the MAP and the respective HDI from the merged posterior distributions. We further checked the Hill stability of the system (Sundman, 1913) when assuming the entire merged posterior distributions by calculating the angular momentum deficit (AMD, Laskar, 1997, 2000; Laskar & Petit, 2017) criterion (Eq. 26 from Petit et al., 2018). Table 5 lists the parameters returned by TRADES (MAP and HDI) along with their respective priors.
@@
After extracting all the synthetic transit timings (i.e. the “observed”: O) from TRADES’ analysis, we computed the “calculated” (C) counterpart according to the linear ephemerides model based on what listed in Table 2. We plotted the as a function of time as well as the RV best-fit model in Figures 6 and 7.
@@
Additionally, we investigated if the best-fit configuration is in or close to a MMR. We integrated the MAP parameters with the N-body code rebound (Rein & Liu, 2012) and the symplectic Wisdom-Holman integrator whfast (Rein & Tamayo, 2015; Wisdom & Holman, 1991) for 10 000 years. We computed the evolution of the critical resonance angles of TOI-396 b and TOI-396 c
@@
where and (for a order 5:3 MMR), while is the longitude of the pericentre. We also computed the evolution of . In case of MMR, we expect that both and librate (i.e. oscillate) around a fixed value for the entire orbital integration. Instead, if these angles circulate, that is they span the full 0∘–360∘ range (or equivalently the – range), then the planet pair is not in resonance. We found that both and circulate (see the two upper panels of Fig. 8), which indicates that the system is not in an exact 5:3 MMR. Even if seems to oscillate around , it circulates every years (bottom panel of Fig. 8), which further confirms that the system is not trapped in an MMR state. We also found the same behaviour for the resonant angles of 200 random samples drawn from the posterior distribution (not shown here). Our conclusions are consistent with the simulations performed by Vanderburg et al. (2019), who found that most realisations of the system are not in resonance.
@@
As shown in Fig. 6, the MAP parameter set predicts that the TTV super-period is longer than the time spanning the two clustered TESS observations. We decided to track the potential evolution of the TTV signals over a temporal baseline of 5.2 years. To this end, we ran forward numerical N-body simulations with TRADES, setting the initial conditions of the orbital parameters to be integrated at . We ran a first simulation taking all the parameters from the MAP solution. Then, we further ran 200 simulations, where the sets of the system’s parameters were randomly drawn from the merged posterior distributions. We computed the synthetic (i.e. observed: ) transit times and created a simulated plot against time, where the counterpart of was computed assuming the linear ephemerides of Table 2. The results are displayed in Fig. 9 that emphasises a progressive drift of the values inferred from the MAP parameters (black line) with respect to the zero value. As a consequence, the TTV amplitudes of planets b and c increase with time. In particular, the semi-amplitude of the black curves are about two and five hours for b and c, respectively. The TTV super-period seems to be roughly equal to or larger than the integration time span, that is 5 years. The variance of the TTV amplitude (shaded area) is inferred from the results of the additional 200 simulations and reflects the widths of the posterior distributions from which the system’s parameters were drawn. The remarkable drifts (i.e. the poorly constrained linear ephemerides) combined with the uncertainty on the TTV amplitudes make challenging to plan future observations of planets b and c, as the actual transit timings might differ from the linear ephemerides predictions by 5 and 10 hours for TOI-396 b and TOI-396 c, respectively. TESS will not observe the target in the foreseeable future.191919https://heasarc.gsfc.nasa.gov/wsgi-scripts/TESS/TESS-point_Web_Tool/TESS-point_Web_Tool/wtv_v2.0.py/

### v2

@@ 5.1 Joint LC and RV analysis with linear ephemerides
As mentioned in Sect. 4, we set and as free parameters under the control of a uniform prior, which implies assuming linear ephemerides. With this setup, we improved the transit depth precision of all three planets by a factor 1.4 if compared with the results of Vanderburg et al. (2019). This improvement level is consistent with having twice the number of data points with respect to the LC analysis performed by Vanderburg et al. (2019), as well as with low TTV amplitudes (see Sect. 5.2).
@@
When combining the mass and radius values of the three planets, we obtain the following median estimates for the bulk planetary densities: , (), and g cm-3. We note that the RV-undetected TOI-396 c would be the least dense planet, while the densest planet is the outermost one (i.e. TOI-396 d), which constitutes a quite atypical architecture within the observed exoplanet population (e.g. Ciardi et al., 2013; Weiss et al., 2018; Mishra et al., 2023). However, this conclusion is just tentative, given the uncertainties on the mean planetary densities. Moreover, the detection level of the RV-inferred parameters of TOI-396 c is not statistically significant. We further tested whether including the MINERVA-Australis (Addison et al., 2019) RV data (30 measurements as taken from Vanderburg et al., 2019) can help detect the elusive planet. However, it turned out that their precision level ( 6 m s-1) is not high enough to improve the characterisation of the system. In other words, the RV semi-amplitudes we obtained are consistent and indistinguishable within the statistical fluctuation with what was derived from the more precise HARPS data set. This led us to further check that TOI-396 c indeed belongs to this system (Sect. 5.2) and to investigate the reason for its RV non-detection (Sect. 5.3).
5.2 Joint LC and RV analysis accounting for TTVs
@@
In line with Vanderburg et al. (2019), we also confirmed that the period ratio is commensurable and differs from the 5:3 ratio by less than 0.027%. As planets with orbits in, or close to, resonances are likely to show TTVs, we decided to repeat the same MCMC analysis outlined above, but enabling the transit timings of each transit event to vary to then compute the TTV amplitude with respect to the linear ephemerides model derived in Sect. 5.1. All jump parameters converged (). The medians of the posterior distributions of the most relevant system parameters along with the 68.3% confidence intervals are listed in Table 2. The phase-folded LCs of the three planets are shown in Figure 1, while the phase-folded RV time series are displayed in Figure 2. In particular, the middle panel of Fig. 2 shows that, after subtracting the RV signals of both planets b and d, the RV time series looks flat consistently with the non-detection of TOI-396 c.
@@
This analysis gives , , and and confirms the RV detection of TOI-396 b and TOI-396 d (at the 3.8 and 4.5-level, respectively) with and . These outcomes are consistent with the results obtained from the analysis that assumes linear ephemerides (Sect. 5.1).
Fig. 3 shows the three planets of the system (star symbol) along with other exoplanets whose density is more significant than the 3 level999The planet properties were downloaded from the Nasa Exoplanet Archive as of 5 June 2024: https://exoplanetarchive.ipac.caltech.edu/ (circular marker) in the mass-radius (MR) diagram. The superimposed theoretical MR models as taken from Aguichine et al. (2021, A21) and Haldemann et al. (2024) help guide the eye; however, we warn the reader of possible degeneracies occurring in the MR plane. A thorough internal structure analysis of the well characterised TOI-396 b and d is presented in Sect. 7.
@@
5.3 Discussion regarding why TOI-396 c is not detected in the RV time series
@@
In order to investigate this hypothesis, we performed a frequency analysis of the line profile variation diagnostics (FWHM, contrast, and skewness) and activity indicators (log R and H indexes). Figure 5 displays the time series (left column), along with the respective generalised Lomb-Scargle (GLS, Zechmeister & Kürster, 2009) periodograms (right column). We assessed the significance of the peaks detected in the power spectra by estimating their false alarm probability (FAP), that is the probability that noise could produce a peak with power higher than the one we found in the time series. To account for the possible presence of non-Gaussian noise in the data, we estimated the FAP using the bootstrap randomisation method (see, e.g., Murdoch et al., 1993; Kuerster et al., 1997; Hatzes, 2019). Briefly, we computed the GLS periodogram of 105 mock time series obtained by randomly shuffling the data points along with their error bars, while keeping the timestamps fixed. We defined the FAP as the fraction of those mock periodograms whose highest power exceeds the power of the real data at any frequency. In the present work, we considered a peak to be significant if its false alarm probability is FAP 0.1 %.
@@
To check whether a quasi-periodic signal compatible with 7 d is also present in the photometric data, for each TESS sector we extracted custom LCs from pixel data using lightkurve (Lightkurve Collaboration et al., 2018). In detail, we adopted the default quality bitmask and set the aperture to ‘all’, which corresponds to an aperture larger than the one used by the official SAP pipeline. In fact, larger apertures mitigate the effect of slow image drifts that could interfere with slow flux changes, such as the 6–8 d rotation period signals we aim to detect. After removing the temporal windows containing the transit events, we computed the GLS periodograms of these lightkurve-based LCs. The FAP was computed following the same bootstrap technique outlined above for the RV activity indicators. The four periodograms (Fig. 15, first column) exhibit very significant peaks at 7.7, 7.9, 7.5, and 6.8 days for TESS Sectors 3, 4, 30, and 31, respectively. Except for Sector 30, they are not the most prominent peaks; however, they persist even after removing the most significant signals (Fig. 15, second column). Whereas we acknowledge that the likely rotation period of the star is close to the first harmonic of the orbital period of TESS around the Earth ( 14 days), the photometric signal at 6–8 d is significant in all the four TESS sectors and persists after pre-whitening the data. This signal is consistent with the rotation period detected in the HARPS activity indicators and inferred from log R, , and gyrochronology, suggesting it is astrophysical in nature and due to the presence of active regions carried around by stellar rotation. Assuming the log R-based d as our reference estimate, the orbital period of TOI-396 c ( d) is close to , which may explain the non-detection of planet c within the RV time series.
@@
As is not exactly equal to , we then repeated the test outlined above, but this time we injected into the original HARPS time series artificial Keplerian signals with . The values obtained by the MCMC analyses depending on the different values are reported in Table 4. The values are systematically and significantly smaller than the corresponding values, which a posteriori supports the conclusion that the stellar rotation period is around 6–8 d. Furthermore, a planet with this orbital period would be firmly detected if its RV semi-amplitude were greater than , which is the largest RV semi-amplitude detected for this system. Instead, by injecting a Keplerian signal with and , the MCMCI analysis is able to barely detect (at 2) a planetary signal whose amplitude is half of that expected. The detection level increases when increases; however, we still underestimate . These tests prove that it is difficult to retrieve planetary signals with .
In summary, we conclude that stellar activity is responsible for generating spurious RV signals whose harmonics also include the stellar rotation period. As a consequence, it is hard to reliably detect planets with orbital periods comparable to via the RV technique and Table 4 quantifies the magnitude of this effect. We note that the we obtained from the MCMCI analysis in Sect. 5.2 is comparable to the retrieved when inserting an artificial signal having m s-1, which let us to speculate that TOI-396 c might have 3.0 and 2.0 g cm-3.
@@ 6 Joint RV and TTV dynamical analysis
As mentioned, the anti-correlation pattern of the observed TTVs (see Fig. 14) is a typical signal of the dynamical interaction between TOI-396 b and c. However, the data in our hands are not enough to currently derive meaningful planetary masses from the TTVs. Indeed, the photometric observations are clustered in two groups (about two years apart), which results in a partial coverage of the curvature of the TTVs. This prevents us from accurately mapping the full phases and super-periods of the TTV signals. Hence a dynamical fit onto the transit times would lead to a high fraction of low-amplitude TTVs with a short super-period or to a low fraction of high-amplitude long-period TTV signals.
@@
Despite this, we attempted to run a dynamical joint fit of the TTV and RV data set with TRADES141414https://github.com/lucaborsato/trades (Borsato et al., 2014, 2019, 2021), a Fortran-python code developed to model TTVs and RVs simultaneously along with N-body integration. We have taken and fixed the stellar mass and radius values from Table 1. We further fixed the orbital inclinations to the values in Table 2 and the longitude of the ascending nodes to 151515 We assumed the same reference system of Winn (2010) and Borsato et al. (2014) and later used in Borsato et al. (2019, 2021, 2024); Nascimbeni et al. (2023, 2024). for all the planets. We fitted the planetary masses scaled by the stellar mass (), the orbital periods (), the eccentricities () and the arguments of the pericentre () in the form (, ), and the mean longitudes (161616The mean longitude is defined as , where is the argument of the pericentre, is the longitude of the ascending node, and is the mean anomaly.). We also fitted for an RV offset () and for an RV jitter term () by adopting as step parameter, although we used the de-trended RV data set as derived from Sect. 5.2, where a jitter term was already included in the RV errors. All fitting parameters were subject to uniform priors that account for their respective physical boundaries (see Table 5); for the eccentricities we applied a log-penalty () based on the half-Gaussian () from Van Eylen et al. (2019).
@@
The reference time for the dynamical integration of the orbital parameters was set at , that is before all available observations. We combined the quasi-global differential evolution (Storn & Price, 1997) optimisation algorithm implemented in PyDE (Parviainen et al., 2016) with the Affine Invariant MCMC Ensemble sampler (Goodman & Weare, 2010) emcee implemented by Foreman-Mackey et al. (2019, 2013). We first ran PyDE and evolved 68 different initial configurations of parameter sets for generations (number of steps for which each parameter is evolved). To perform the dynamical analysis in an MCMC fashion, we then ran emcee, assuming as starting point the outcome obtained with PyDE. We set up 68 chains for steps each. To sample the parameter space efficiently, we mixed the DEMove() and DESnookerMove() differential evolution moves171717See emcee documentation: https://emcee.readthedocs.io/en/stable/tutorials/moves/ in the proportion 80%–20% (Nelson et al., 2014; ter Braak & Vrugt, 2008). We repeated the sequence PyDE+emcee twice, with different seeds for the random number generator. The chains reached convergence according to visual inspection and statistical indicators, such as the Gelman-Rubin , the Geweke’s statistic (Geweke, 1991), and the auto-correlation function181818See e.g. Goodman & Weare (2010) and the emcee documentation at https://emcee.readthedocs.io/en/stable/tutorials/autocorr/#autocorr. For both runs, we applied a conservative thinning factor of 100 and discarded the first of the chains (burn-in). For each run, we derived the reference outcome (hereinafter also referred to as best-fit), as the maximum-a-posteriori (MAP) parameters’ set, that is the set of parameters that maximises the log-probability191919 The log-probability is given by , where is the log-likelihood and is the log-penalty of the eccentricities. . The uncertainty of each parameter is quantified by the high-density interval (HDI) at 68.27%202020 An HDI at 68.27% is the equivalent of the percentile range of a Normal distribution that provides the 1 standard confidence interval. of its posterior distribution. For each parameter we computed a Z-score defined as , where the subscripts denote the two different runs and . It turned out that for each parameter, which allows us to merge the posterior distributions deriving from the two runs to finally compute the MAP and the respective HDI from the merged posterior distributions. We further checked the Hill stability of the system (Sundman, 1913) when assuming the entire merged posterior distributions by calculating the angular momentum deficit (AMD, Laskar, 1997, 2000; Laskar & Petit, 2017) criterion (Eq. 26 from Petit et al., 2018). Table 5 lists the parameters returned by TRADES (MAP and HDI) along with their respective priors.
@@
After extracting all the synthetic transit timings (i.e. the ‘observed’: O) from TRADES’ analysis, we computed the ‘calculated’ (C) counterpart according to the linear ephemerides model based on what listed in Table 2. We plotted the as a function of time as well as the RV best-fit model in Figures 6 and 7.
@@
Additionally, we investigated if the best-fit configuration is in or close to an MMR. We integrated the MAP parameters with the N-body code rebound (Rein & Liu, 2012) and the symplectic Wisdom-Holman integrator whfast (Rein & Tamayo, 2015; Wisdom & Holman, 1991) for 10 000 years. We computed the evolution of the critical resonance angles of TOI-396 b and TOI-396 c
@@
where and (for a second order 5:3 MMR), while is the longitude of the pericentre. We also computed the evolution of . In case of MMR, we expect that both and librate (i.e. oscillate) around a fixed value for the entire orbital integration. Instead, if these angles circulate, that is they span the full 0∘–360∘ range (or equivalently the – range), then the planet pair is not in resonance. We found that both and circulate (see the two upper panels of Fig. 8), which indicates that the system is not in an exact 5:3 MMR. Even if seems to oscillate around , it circulates every years (bottom panel of Fig. 8), which further confirms that the system is not trapped in an MMR state. We also found the same behaviour for the resonant angles of 200 random samples drawn from the posterior distribution (not shown here). Our conclusions are consistent with the simulations performed by Vanderburg et al. (2019), who found that most realisations of the system are not in resonance.
@@
As shown in Fig. 6, the MAP parameter set predicts that the TTV super-period is longer than the time spanning the two clustered TESS observations. We decided to track the potential evolution of the TTV signals over a temporal baseline of 5.2 years. To this end, we ran forward numerical N-body simulations with TRADES, setting the initial conditions of the orbital parameters to be integrated at . We ran a first simulation taking all the parameters from the MAP solution. Then, we further ran 200 simulations, where the sets of the system’s parameters were randomly drawn from the merged posterior distributions. We computed the synthetic (i.e. observed: ) transit times and created a simulated plot against time, where the counterpart of was computed assuming the linear ephemerides of Table 2. The results are displayed in Fig. 9 that emphasises a progressive drift of the values inferred from the MAP parameters (black line) with respect to the zero value. As a consequence, the TTV amplitudes of planets b and c increase with time. In particular, the semi-amplitude of the black curves are about two and five hours for b and c, respectively. The TTV super-period seems to be roughly equal to or larger than the integration time span, that is 5 years. The variance of the TTV amplitude (shaded area) is inferred from the results of the additional 200 simulations and reflects the widths of the posterior distributions from which the system’s parameters were drawn. The remarkable drifts (i.e. the poorly constrained linear ephemerides) combined with the uncertainty on the TTV amplitudes make challenging to plan future observations of planets b and c, as the actual transit timings might differ from the linear ephemerides predictions by 5 and 10 hours for TOI-396 b and TOI-396 c, respectively. TESS will not observe the target in the foreseeable future.212121https://heasarc.gsfc.nasa.gov/wsgi-scripts/TESS/TESS-point_Web_Tool/TESS-point_Web_Tool/wtv_v2.0.py/